OpenStudy (anonymous):

x^2+y^2+4x+6y-40=0 x-y=10

OpenStudy (anonymous):

i need solution

OpenStudy (anonymous):

x=5, y = -5 and x=0, y=-10

OpenStudy (anonymous):

solution

OpenStudy (kira_yamato):

\[x^2+y^2+4x+6y-40=0 ---(1)\] \[x = y + 10---(2)\] Sub (2) into (1): \[(y+10)^2+y^2+4(y+10)+6y-40=0\] \[y^2+20y+100+y^2+4y+40+6y-40=0\] \[2y^2+30y+100=0\] \[(y+5)(y+10)=0\] \[y = -5 OR y = -10\] Substitute y = -5 \[x = -5+10 = 5\] Substitute y = -10 \[x = -10+10 = 0\] So, x = 5, y = -5 OR x = 0, y = -10